Institut for Matematik og Datalogi Odense Universitet April 19, 1996

JFB

DM11 – F96 – Lecture 11

Lecture, April 17

We finished Appendix 3 and covered sections 5.3, 6.1, and up through definition 2 of section 6.2.

Lecture, April 24

We will continue in chapter 6.

Lecture, May 1

We will finish chapter 6, skipping topological sort, which is covered in DM02. We will also begin on Huffman codes.

Announcement

On Thursday, April 25, the discussion section D2 from U13 will be held in U22.

Problems to be discussed in week 18

• 24 and 25 from section 5.3.

• 25, 27, 28, and 33 from section 6.1. Note that 28c uses the symmetric difference, which is defined on page 56.

• 7 and 11 from section 6.2.

• 4, 9, and 11 from section 6.3.

• 12, 14, 26, 28, and 32 from section 6.4.